OpenStudy (anonymous):

find the sum of the series ∑ (x^n+1)/(n+1)

OpenStudy (anonymous):

help me

OpenStudy (shubhamsrg):

n from 0 to inf ? And also what is x ?

OpenStudy (anonymous):

is that\[\sum_{n=0}^\infty \frac{x^{n+1}}{n+1}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure about your question? what do they want? finding sum of series or convergence radius or something?

OpenStudy (anonymous):

yes...very sure....

OpenStudy (shubhamsrg):

So its like x + x^2 /2 + x^3 /3 .... Looks like taylor expansion for e^x . Just the slightest of manipulations

OpenStudy (anonymous):

well \(e^x\) series is\[\sum_{n=0}^\infty \frac{x^{n}}{n!}\]

OpenStudy (shubhamsrg):

Oh lol i am sorry :P

OpenStudy (anonymous):

ratio test will give you the interval \(|x|<1\) as convergence radius but I'm not sure about the sum in that interval, let's work on it :-)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

@RadEn @ganeshie8 idea?

OpenStudy (anonymous):

oh no, wait a min

OpenStudy (anonymous):

if \(|x|<1\) you can write\[\frac{1}{1-x}=1+x+x^2+...\]take integral of both sides

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

@ferwanz am I clear? :-) tnx @ganeshie8

OpenStudy (anonymous):

yep.... tq everyone...